Year (x) 2005 2006 2007 2008 2009 2010 2011 2012 Population (y, in million) 85.26 86.97 88.71 90.46 91.02 92.6 94.18 95.77 Source: tradingeconomic.com.philippines/population 1. Find the regression line that will predict the population for any given year.


1
Expert's answer
2021-05-24T12:37:35-0400

Using online calculator www.graphpad.com, we get an equation of the regression line:

"y=1.451x-2823"



Calculations:

"y=ax+b"


"a=\frac{n\sum xy-\sum x\sum y}{n\sum x^2-(\sum x)^2}"


"b=\frac{\sum y\sum x^2-\sum x\sum xy}{n\sum x^2-(\sum x)^2}"